Tech Staff Meeting Summary 6/3/02

Current Top Priorities are:Complete DBB, ANSS

Reorganize this list to get it back into some more clear priority order.

Greg was ruminating and came up with this todo list. Much of it that relates to other staff has been incorporated here. Til I figure out what to do with the rest of it, it is in unedited form.

computer room. Mitch, Steve, and Chris W. will move tables and benches around to make space for the worm racks. Need to come up with a plan for the other room.

ANSS. Permitting is going nicely. Received 9 of 20 CMG-5TD's. One is undergoing testing. Mitch is testing software (gcf2ew, new scream2ew, new scream). Greg is working on a new installation design that should be easier to install. We will be swapping the NSMP our k2's from last year with their CMG5TD for this year. Arnie Acosta will assist.

CVTN is running. Need one more visit to get comm going.

CBHT comm. Switch combined with dramatic k2ew config changes appeared to work. Steve said a new firmware upgrade may also solve the problem.

HDBT. Need a fence and gravel for ground cover.

NAIT. NWS does not know when the VPN will be installed. Mitch will ask the NWS Houston contact for more info.

CUET. Provided additional information to Pat Bohannan at TEMA regarding the software connections. He's working on the firewall issues (we hope).

SFTN. K2 and MSS100 are fine. No comm out the ethernet. Needs debugging.

TUMT. Mitch reconfigured all k2ew modules to be similar to CBHT (i.e. more robust with dropouts). Seems to have fixed TUMT. Now it seems to run out of open connections and only a reboot will fix (since you can't telnet in if there are no connections). Need to look into why sessions don't close.

DBB. Steve suggested a terminal server/watchdog to do dynamic resets when they hang. Has been working great though so this is backburner (need to order a bigger backburner and more pilot lights). ISTI worm module to monitor outages should best wait until RPC is running. Jim will check into getting the dish hung at NMAD and LNXT.

NMAD fence is done.

GPS. Got a repaired receiver back from UNAVCO. It needs to be installed at Troy. The phone line at CVTN is noisy causing the download to time out due to resulting reduced bandwidth. Marcello will need to manually download the data and work on a wireless solution to get internet going. McCarthy lantronix box needs a reset.

MPHB and the box in the yard is up--sort of. Redirector software has been uninstalled. Mitch is working with testing several packages mentioned above under ANSS and they seem to be working. Should have MPHB online soon.

ET. Mitch shut down the lrnc heli-pages. John is focussing on new installs in east TN. The plan is to completely rebuild the NC stations when we can get time to get over and help. Then we'll put lrnc back online. Receiving all USNSN stations in Memphis. Steve replaced the computers for athn. Now there are no more bsd boxes. Worms running both in the EMA and on star mountain and streaming data. Still need to work out the details of getting data out of athn. Don't have the bandwidth for waveforms, but might send statrig messages and form trigger groups in Memphis. John found a bad preamp in in CPRT and is repairing. Repeater at FGTN is installed.

NM. HALT has a timing problem. Not sure specifically what other stations need work but we're in pretty good shape. Need to work in a routine maintenance trip this summer (e.g. for knocking down the kudzu, checking voltages, etc).

Chapman and Munsey are working on a site for Oakridge to put a CMG3 or 40T. Mitch will provide more info on proposed KY NSN station to get contacts.

FDL. Have recorded several events on the NY stations. Will be acquiring 2 new k2's and a spare episensor deck plus 4 new CMG-6TD systems. New shipping containers went out for bid. Mitch is haggling with Bodin on how best to run this lab. Clearly maintenance needs to be addressed but it is not clear that we can have one person know everything about all the equipment.

NHIN. Steve is monitoring. AASG meeting is June 22 to 26.

PII. Most of PI has been cleaned out. PII proposal for New Madrid was not funded.

Monroe. Still want to. But when and where?

RPC. Parts are being ordered. Sheila set up a constructed equipment account. Mitch is keeping a folder.

FY02 ANSS. Will be installing 20 CMG5TD systems with vaults and accessories as close to last year as possible. See the ANSS page for station details.

I-40 Bridge. Progressing nicely. Can probably do some installations later this year or early next year.

Need to do site visits of Lamont stations sometime this year.

3904/215E. Will expand into the closet and be used as a computer setup/testing lab.

Generator. Specs are getting approved by P-plant after which it will go out for bid. Hopfully we can get it awarded this month.

Purchase Parker Vehicle. Done.

Giddings rebuild parts have been ordered. Randy is drilling some cores Monday, June 10. Chris will accompany him on the first one in Shelby Forrest to provide some instruction. Randy is trying to get some other interested parties there.

ESEE. Postponed to October. probably have a bunch of refteks and streckeisens. This may not work because the sensors are too temperature sensitive.

Auburn. Wants to make some cores. They submitted a proposal. If funded will need our help getting a couple cores. Shouldn't be a big deal. A day to two maybe.

Repair former SWET CMG-3.

Need permission for additional equipment at Autozone.

New ANSS vault design. Need to start ordering part. Need a design to budget to get the money to start ordering parts.

Mitch will contact Dan Kohr regarding P-Plant plans for the Liebert.

EW racks are ready to assemble. Or at least one is, need to get the other two from Wessco.

Install air compressor.

Calibrate S-13's.

SFTN ethernet redesign.

Clean out FDL attic.

Battery charging station

renew 219 band licenses.

reattach 3904 gnd straps.

Remove VSAT. done.

Repair fried SFTN k2 panel.